the measure of an angle is 12 degrees less than twice the measure of its supplement. what is the measure of the angle?
Answer by checkley77(12844)

x+y=180 degrees.
x=2y-12
(2y-12)+y=180
3y-12=180
3y=180+12
3y=192
y=192/3
y=64 degrees.
180-64=116 degrees.
Proof:
64+116=180
180=180